At 1000C, the vapour pressure of a solution of 6.5 g. of a solute in 100g. water is 732 mm. What is the boiling point of the solution? (Kb = 0.52)

Grade:12

1 Answers

Vikas TU
4 years ago
Raoult's aw of Partial pressure 
Pa^0 - Ps / Ps = nb /na 
Putting the values 
28 / 732 = 6.5 * 18 / Mb *100 
Mb = 30.6 
Del Tb = 0.52 * 6.5 * 1000/ 30.6 * 100 
=> 1.10 
Boiling point = 100+ 1.10 
=> apx 101 degree centigrade